OpenStudy (samigupta8):

prove that the vectors A=3i+2j-2k,B=-1+3j+4k and c=4i-j-6k can form a triangle

OpenStudy (goformit100):

you can too solve it using Martices.

OpenStudy (badhi):

take \(\vec A -\vec B\). you'll get \[(3 \mathbb i +2\mathbb j-2\mathbb k)-(-\mathbb i+3\mathbb j+4 \mathbb k)=4\mathbb i - \mathbb j -6\mathbb k=\vec C\] so eventually you get \(\vec A -\vec B=\vec C\) or \[\vec A=\vec B +\vec C\] This a property when three vectors form a triangle

OpenStudy (goformit100):

Using Matrix will be much easier